OpenStudy (doctor):

A sample of gas has a pressure of 3.2 atmospheres at a temperature of 301 Kelvin. What would be the pressure of the same sample of gas when the temperature is decreased to 285 Kelvin? Show all of the work used to solve this problem.

sam (.sam.):

Gay-Lussac's law \[\frac{P_1}{T_1}=\frac{P_2}{T_2}\] \[P_2=\frac{P_1T_2}{T_1}\] Plug and chug

OpenStudy (doctor):

THANK YOU SOOOOO MUCH but do you think I could ask you one more question?